Wallis Company manufactures only one product and uses a standard cost system. The company uses a predetermined plantwide overDuring the year Wallis completed the following transactions: a. Purchased (with cash) 237,500 pounds of raw material at a priReq 4 Req 1 Req 2 and 3 Compute all direct materials, direct labor, and fixed overhead variances for the year. (Indicate theRea 2 and 3 Req 1 Req 4 Record transactions a through I for Wallis Company. Compute the ending balances for Wallis CompanysReq 2 and 3 Req 4 Req 1 Prepare Wallis Companys income statement for the year. (Enter your dollars in thousands. Round your
Answer #1

Part 1

Materials price variance = AQ*(AC-SC) = 237500*(31-33) = $475000 F

Materials quantity variance = SC*(AQ-SQ) = 33*(218750-(96500*2)) = $849750 U

Labor rate variance = AH*(AR-SR) = 248000*(16-15) = $248000 U

Labor efficiency variance = SR*(AH-SH) = 15*(248000-(96500*3)) = $622500 F

Budget variance = actual FOH – Budgeted FOH = 2747500-2899000 = 151500 F

Volume variance = Budgeted FOH – Applied FOH = 2899000-(96500*3*10) = 4000 U

Part 2 & 3

Wallis company

Transaction analysis

For the year ended 12/31/XX

(dollars in thousands)

Cash

Raw materials

Work-in-progress

Finished goods

PP&E (net)

=

Material price variance

Material quantity variance

Labor rate variance

Labor efficiency variance

Fixed overhead budget variance

Fixed overhead volume variance

Retainer earnings

1/1

850000

300000

420000

10000000

11570000

a.

(7362500)

7837500

475000

b.

(7218750)

6369000

(849750)

c.

(3968000)

4342500

(248000)

622500

d.

(1355000)

2895000

(1392500)

151500

(4000)

e.

(13606500)

13606500

f.

15895000

15895000

g.

(13183500)

(13183500)

h.

(2127500)

(2127500)

i.

(475000)

849750

248000

(6225000)

(151500)

4000

147250

1932000

918750

0

843000

8607500

0

0

0

0

0

731250

Part 4

Income statement

Sales

15895000

Cost of goods sold (standard)

13183500

Total variance adjustments

(147250)

Cost of goods sold (actual)

13036250

Gross profit (actual)

2858750

Selling and administrative expenses

2127500

Net income

$731250